how much more is 2/5 than 1/7 in mixed fraction?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

how much more is 2/5 than 1/7 in mixed fraction?
---------
2/5 - 1/7
-----
Find a common denominator, subtract

Answer by stanbon(75887)

how much more is 2/5 than 1/7 in mixed fraction?
=====
(2/5)-(1/7) = (7*2-5*1)/35 = 9/35 = 0.2571
============
Cheers,
Stan H.